OpenStudy (anonymous):

checkmy worrrrk?(: What is the simplified form of 2 over x squared minus x minus 1 over x ? x minus 1 over the quantity x times x plus 1 1 minus x over the quantity x times x plus 1 3 minus x over the quantity x times x minus 1 x plus 2 over the quantity x times x minus 1 I got A...

zepdrix (zepdrix):

Your sentences are difficult to read sometimes :D lol Is it suppose to look like this?\[\large \frac{2}{x^2}-x-\frac{1}{x}\]

OpenStudy (anonymous):

te x is supposed to be with the x^2 under 2..sorry Idk how to paste it here without it turning this way/;

zepdrix (zepdrix):

\[\large \frac{2}{x^2-x}-\frac{1}{x}\] Ah ok :O i see. sec, working through it.

OpenStudy (anonymous):

okay!

zepdrix (zepdrix):

Hmm, I don't think it's A. Want to see some of the steps to figure out where u went wrong? :U

OpenStudy (anonymous):

yes!

zepdrix (zepdrix):

We need a common denominator, which in this case will be the product of the denominators. \(\large x(x^2-x)\) So it looks like our first fraction needs to be multiplied by \(\large \cfrac{x}{x}\) while our second fraction should be multiplied by \(\large \cfrac{x^2-x}{x^2-x}\) Which will give us,\[\large \frac{2x}{x(x^2-x)}-\frac{1(x^2-x)}{x(x^2-x)} \qquad=\qquad \frac{2x-(x^2-x)}{x(x^2-x)}\]

zepdrix (zepdrix):

Lemme know if any of those steps are too confusing :o

zepdrix (zepdrix):

That's how we would combine the fractions, then from there we can simplify by doing some fancy stuff.

OpenStudy (anonymous):

I got that

zepdrix (zepdrix):

Distributing the negative on top gives us,\[\large \frac{2x-x^2+x}{x(x^2-x)} \qquad=\qquad \frac{3x-x^2}{x(x^2-x)}\]Did you remember to distribute that negative sign? :)

OpenStudy (anonymous):

ohhh no. I didn't do that

zepdrix (zepdrix):

Bahh easy spot to make a mistake D: poor gal.

OpenStudy (anonymous):

so is it C?

zepdrix (zepdrix):

Yay good job \c:/

OpenStudy (anonymous):

thank you so much! big help!(: